# Feature Attribution with Necessity and Sufficiency via Dual-stage Perturbation Test for Causal Explanation

### v1

@@ 6.1 Experimental Setup
1) Datasets and models. We utilized six public datasets with well-trained models: CIFAR10 (Krizhevsky et al., 2009) with ResNet18, MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017) with LeNet5, as well as BACommunity, Pubmed, and Citeseer with GCN (Ying et al., 2019; Sen et al., 2008; Kipf & Welling, 2016). CIFAR10, MNIST, and Fashion-MNIST are image datasets. BACommunity, Pubmed, and Citeseer are graph datasets. Please see Appendix B.1 for details.
[…]
3) Evaluation Metrics. We utilized six widely used metrics to evaluate the effectiveness of FANS: Infidelity(INF) (Yeh et al., 2019), Iterative Removal Of Features(IR) (Rieger & Hansen, 2020), ${}^{+}$(${}^{+}$), ${}^{-}$(${}^{-}$) (Amara et al., 2022), Max-Sensitivity(MS) (Yeh et al., 2019), Sparseness(SPA) (Chalasani et al., 2020), and Recall@12 (Wang et al., 2021) metrics. INF, IR, ${}^{+}$, and ${}^{-}$ measure the extent to which the explanation follows the predicted behavior of the model (i.e., faithfulness), while MS measures the stability of the explanation when subjected to slight input perturbations (i.e., robustness) (Hedström et al., 2023). SPA measures the sparsity of the explanation by using the Gini Index. Recall@$12$ measures how many of the top 12 most important features for a given attribution are features of the “ground-truth explanation”. Remarkably, Recall@$12$ is only suitable for BACommunity since this dataset is synthetic and contains “ground-truth explanation”. Please see Appendix B.3 for details.
To test the explanation methods, we randomly selected 4,000 images from each image dataset (i.e., MNIST, Fashion-MNIST, and CIFAR10) and adopted the same configuration as the literature (Wang et al., 2021) on the graph datasets (i.e., BACommunity, Citeseer, and Pubmed). The results presented in Tables 1 and 2 are largely averages of the metric values for the different inputs, with the exception of ${}^{+}$ and ${}^{-}$, since these are metrics designed for the entire dataset rather than an individual sample. As each graph dataset we use contains only one graph, and FANS requires multiple samples for SIR, we create a graph-augmented dataset for each graph by randomly masking their adjacency matrix. More implementation details are presented in Appendix B.4.
@@ 6.2.1 Quantitative evaluation
Robustness analysis. Next, we investigate the robustness of FANS and the baseline methods by examining the effect of introducing noise. Following the literature(Yeh et al., 2019), we designed a robustness experiment as follows. Given various values of $r$, for $r\in[0,1]$, we add noise to each pixel of input image $\mathbf{x}_{i}$ from the interval [$-r$, $r$]. We then evaluate the difference between the explanations of each perturbed image and its original explanation via the Max Sensitivity (MS) metric. As shown in Fig.3(a)(b), as expected, all methods demonstrate a gradual deterioration in performance with increased perturbation strength. However, the slope of this deterioration is considerably lower using FANS, which demonstrates far superior robustness on both examined datasets. This can be explained by the fact that FANS learns from multiple samples which are similar to the target input through SIR, rather than a single sample, which renders it more robust to perturbation noise.
Figure 3: Performance on robustness comparison in image datasets CIFAR10 and MNIST under different perturbation strengths.
(a) Citeseer
(b) Pubmed
Figure 4: Ablation study of sufficiency (SF) module, necessity (NC) module, and SIR-based Abduction (AD) on graph datasets Citeseer and Pubmed.
6.2.2 Qualitative evaluation
Figure 5: Attributions visualization on the MNIST dataset.
6.3 Ablation study
@@
The core components of FANS are the necessity, sufficiency, and SIR-based Resampling modules. In order to validate the importance of each component and better understand their relationship to different explainability metrics, we ablate our method by iteratively removing each component. We refer to FANS without the sufficiency module as FANSw/oSF, FANS without the necessity module as FANSw/oNC, and FANS without SIR-based Resampling as FANSw/oSR. We visualize the effects of this ablation on performance relative to the full algorithm in Figure 4 for the Citeseer and Pubmed datasets. In Figure 4(a)(b), we first remark that the full FANS algorithm is the only version of the method to perform uniformly well across both datasets and all metrics. When comparing with FANSw/oSF, we notice a slight decrease in the faithfulness metric FID+ and a significant decrease in FID-, while the sparsity metric remains stable, where FID- and FID+ evaluate the proportions of predictions remain unchanged and change when features with high attribution scores are retained and removed, respectively. Conversely, FANSw/oNC exhibits an opposite trend to the above three metrics of FANSw/oSF, with FID+ becoming significantly worse, FID- remaining stable, and SPA becoming worse. This result is expected since the sufficiency and necessity modules ensure that the feature subset contains all features responsible for the prediction and removes those that are not respectively, which result in a more dense or sparse optimal feature subset. Thus, the optimal subset estimated by FANSw/oSF may only contain a subset of features that are responsible for prediction. Removing these features may lead to prediction change (high FID+), while retaining these features may also lead to prediction change (high FID-) since the important features are incomplete. Similarly, the optimal subset estimated by FANSw/oNC may contain redundancy. Retaining these features may keep the prediction unchanged (low FID-) while removing these features the change in prediction is uncertain (low FID+) due to the redundancy.
@@
Next, we compared FANS with FANSw/oSR. As shown in Fig.4(a)(b), the sparsity metric SPA of FANSw/oSR remains stable because of FANS’s necessity module. However, both faithfulness metrics FID+ and FID- decrease. This may be because FANSw/oSR lacks information about predictions from samples similar to the target input when learning the optimal subset of features, which may lead to FANS being affected by unimportant features in the target input. Additional efficiency analyses involving different methods on CIFAR10 are presented in Appendix 7.
7 Efficiency Analysis
Figure 6: Efficiency analysis on the CIFAR10 dataset.
We evaluate the efficiency of FANS on CIFAR10, in terms of the inference time. We select representative baselines, including optimization-based methods MP, LIME, and CIMI, as well as non-optimization-based methods FeatAblation, Occlusion, DeepLift, and IG.
As shown in Figure 6, the efficiency of FeatAblation, Occlusion, DeepLift, and IG remains stable, as expected since they simply compute gradients or perturbations, enabling parallel computation. However, it is worth noting that when explaining a single sample, FANS is competitive. In contrast, optimization-based methods such as FANS, MP, and LIME tend to have longer inference times, which is expected since they build each optimization instance when explaining each sample. However, FANS is more efficient compared to MP and LIME, probably due to its well-defined objective function without ambiguous trade-off coefficients. Additionally, CIMI adopts mini-batch optimization technology, which further improves its efficiency.
@@ 8 Conclusion
This paper presents a Feature Attribution with necessity and sufficiency (FANS) method an explainability method that can better distinguish the contribution of different features to predictions. Compared with the existing methods, this paper first proposes a general perturbative prediction process of ML models. Building upon this process, we incorporate Probability of Necessity and Sufficiency (PNS) into the process of measuring the contribution of features to prediction through perturbations via our proposed dual-stage perturbation test, which follows the Abduction-Action-Prediction counterfactual reasoning paradigm. To generate counterfactual samples, this paper proposes to use a sampling-importance-resampling approach to approximate the required conditional distribution with the observed samples. Experiment results on six mainstream benchmark datasets further evaluate the effectiveness of our proposed FANS. In summary, this paper takes a meaningful step towards machine learning explainability.
9 Impact Statement
@@
This paper proposes a feature attribution method called FANS, which can distinguish the contributions of different features even if the results change the same after perturbation. FANS can be used in medical diagnosis, autonomous driving, and other fields to assist people in understanding the predictions of the model, which have the potential to facilitate the application of existing artificial intelligence technologies in real-world scenarios.
Appendix A Resampling Weight for Approximate Sampling
@@
SIR consists of three steps. 1) Draw samples $\mathbf{v}_{1}$, …, $\mathbf{v}_{k}$ from the proposal distribution $Q(\mathbf{X})$. 2) For each $\mathbf{v}_{i}$, calculate the weight $w(\mathbf{v})=\frac{P(\mathbf{v}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})\leq c)}{Q(\mathbf{v})}$ if $\|\mathbf{v}_{\mathbf{s}}-\mathbf{x}_{\mathbf{s}}\|_{p}\leq b$ (definition of variable $\tilde{\mathbf{X}}$ in Eq.1) and 00 otherwise. 3) Draw samples from $\mathbf{v}_{1}$, …, $\mathbf{v}_{k}$ based on the their weights.
Next, we propose a method for estimating $\frac{P(\mathbf{v}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})}{Q(\mathbf{v})}$. As the training set of the ML model comes from $P(\mathbf{X})$, we set $Q(\mathbf{X})$:=$P(\mathbf{X})$, thus $w(\mathbf{v})$ =$\frac{P(\mathbf{v}|\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})\leq c)}{P(\mathbf{v})}$ =$\frac{P(\mathbf{v})P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}}|\mathbf{v})}{P(\mathbf{v})P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})}$ =$\frac{P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})|\mathbf{v})}{P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}})\leq c)}$ =$r\cdot P(\bar{A}_{\mathbf{s}},\bar{B}_{\mathbf{s}}|\mathbf{v})$, where $r$ is a normalization constant.
Appendix B Experimental Setup
B.1 Datasets and Target Models

### v3

@@ 5.1 Experimental Setup
1) Datasets and models. We utilized six public datasets with well-trained models: CIFAR10 (Krizhevsky et al., 2009) with ResNet18, MNIST (LeCun et al., 1998) and Fashion-MNIST (Xiao et al., 2017) with LeNet5, as well as BACommunity, Pubmed, and Citeseer with GCN (Ying et al., 2019; Sen et al., 2008; Kipf & Welling, 2016). CIFAR10, MNIST, and Fashion-MNIST are image datasets. BACommunity, Pubmed, and Citeseer are graph datasets. Please see Appendix C.1 for details.
2) Baselines. We selected the most popular or the most recent methods as the baseline. We then compared FANS with six feature-wise attribution methods: Saliency (Simonyan et al., 2013), GuidedBP (Springenberg et al., 2015), IntegratedGrad(IG) (Sundararajan et al., 2017), GradShap (Lundberg & Lee, 2017), DeepLift (Ancona et al., 2018), and IDGI (Yang et al., 2023), as well as twelve feature subset-wise attribution approaches: Occlusion (Zeiler & Fergus, 2014), LIME (Ribeiro et al., 2016), MeaningfulPerturb(MP) (Fong & Vedaldi, 2017), GNNExplainer (Ying et al., 2019), PGExplainer (Luo et al., 2020), FeatAblation (Kokhlikyan et al., 2020), ReFine (Wang et al., 2021), CFGNExplainer (Lucic et al., 2022), CF2 (Tan et al., 2022), NSEG (Cai et al., 2022), CIMI (Wu et al., 2023) and MixUpExplainer (Zhang et al., 2023). The data categories used to test each baseline are largely aligned with their literature. Except that we adjust CIMI to accommodate image data, and run Saliency, GuidedBP, and IG on both graph and image data. Please see Appendix C.2 for details.
3) Evaluation metrics. We utilized six widely used metrics to evaluate the effectiveness of FANS: Infidelity(INF) (Yeh et al., 2019), Iterative Removal Of Features(IR) (Rieger & Hansen, 2020), Fidelity+(FID+), Fidelity-(FID-) (Amara et al., 2022), Max-Sensitivity(MS) (Yeh et al., 2019), Sparseness(SPA) (Chalasani et al., 2020), and Recall@12 (Wang et al., 2021) metrics. INF, IR, FID+, and FID- measure the extent to which the explanation follows the predicted behavior of the model (i.e., faithfulness), while MS measures the stability of the explanation when subjected to slight input perturbations (i.e., robustness) (Hedström et al., 2023). SPA measures the sparsity of the explanation by using the Gini Index. Recall@$12$ measures the number of the top 12 most important features for a given attribution are features of the “ground-truth explanation”. Remarkably, Recall@$12$ is only suitable for BACommunity since this dataset contains “ground-truth explanation”. See Appendix C.3 for details.
To test the explanation methods, we randomly selected 4,000 images from each image dataset (i.e., MNIST, Fashion-MNIST, and CIFAR10) and adopted the same configuration as the literature (Wang et al., 2021) on the graph datasets (i.e., BACommunity, Citeseer, and Pubmed). The results presented in Tables 1 and 2 are largely averages of the metric values for the different inputs, with the exception of FID+ and FID-, since these are metrics designed for the entire dataset rather than an individual sample. As each graph dataset we use contains only one graph, and FANS requires multiple samples for SIR, we create a graph-augmented dataset for each graph by randomly masking their adjacency matrix. See Appendix C.4 for more implementation details.
@@ 5.2.1 Quantitative evaluation
Figure 3: Performance on robustness comparison in image datasets CIFAR10 and MNIST under different strengths of noise $r$.
Robustness analysis. Next, we investigate the robustness of FANS and the baseline methods by examining the effect of introducing noise. Following the literature (Yeh et al., 2019), we designed a robustness experiment as follows. Given various values of $r$, for $r\in[0,1]$, we add noise to each pixel of input image $\mathbf{x}_{i}$ from the interval [$-r$, $r$]. We then evaluate the difference between the explanations of each perturbed image and its original explanation via the Max Sensitivity (MS) metric. As shown in Fig.3(a)(b), as expected, all methods demonstrate a gradual deterioration in performance with increased noise strength. However, the slope of this deterioration is considerably lower using FANS, which demonstrates far superior robustness on both examined datasets. This can be explained by the fact that FANS learns from multiple samples which are similar to the target input through SIR, rather than a single sample, which renders it more robust to perturbation noise.
5.2.2 Qualitative evaluation
Figure 4: Attributions visualization on the MNIST dataset.
5.3 Ablation Study
(a) Citeseer
(b) Pubmed
Figure 5: Ablation study of sufficiency (SF) module, necessity (NC) module, and SIR-based Sampling (SR) on graph datasets Citeseer and Pubmed.
@@ 5.3 Ablation Study
The core components of FANS are the necessity, sufficiency, and SIR-based Sampling modules. In order to validate the importance of each component and better understand their relationship to different explainability metrics, we ablate our method by iteratively removing each component. We refer to FANS without the sufficiency module as FANSw/oSF, FANS without the necessity module as FANSw/oNC, and FANS without SIR-based Sampling as FANSw/oSR. We visualize the effects of this ablation on performance relative to the full algorithm in Figure 5 for the Citeseer and Pubmed datasets. In Figure 5(a)(b), we first remark that the full FANS algorithm is the only version of the method to perform uniformly well across both datasets and all metrics. When comparing with FANSw/oSF, we notice a slight decrease in the faithfulness metric FID+ and a significant decrease in FID-, while the sparsity metric remains stable, where FID- and FID+ evaluate the proportions of predictions remain unchanged and change when features with high attribution scores are retained and removed, respectively. Conversely, FANSw/oNC exhibits an opposite trend to the above three metrics of FANSw/oSF, with FID+ becoming significantly worse, FID- remaining stable, and SPA becoming worse. This result is expected since the sufficiency and necessity modules ensure that the feature subset contains all features responsible for the prediction and removes those that are not respectively, which result in a more dense or sparse optimal feature subset. Thus, the optimal subset estimated by FANSw/oSF may only contain a subset of features that are responsible for prediction. Removing these features may lead to prediction change (high FID+), while retaining these features may also lead to prediction change (high FID-) since the important features are incomplete. Similarly, the optimal subset estimated by FANSw/oNC may contain redundancy. Retaining these features may keep the prediction unchanged (low FID-) while removing these features the change in prediction is uncertain (low FID+) due to the redundancy.
@@
Next, we compared FANS with FANSw/oSR. As shown in Figure 5(a)(b), the sparsity metric SPA of FANSw/oSR remains stable because of FANS’s necessity module. However, both faithfulness metrics FID+ and FID- decrease. This may be because FANSw/oSR lacks information regarding samples that are similar to the target input when learning the optimal subset of features, which may lead to FANS being affected by unimportant features of the target input.
5.4 Convergence Analysis
We corroborate the convergence analysis of FANS under different sizes of $\mathcal{E}$ which are used for SIR-based resampling and show our FANS approach a stationary point. Figure 6 shows the learning process of FANS with sample size $|\mathcal{E}|=20,|\mathcal{E}|=200$ and $|\mathcal{E}|=2,000$. These values correspond to varying degrees of sparsity within the sample set $\mathcal{E}$. The training losses decrease with the number of iterations, leading to a convergent result in all cases. As shown in Figure 6 (a), we find that the training loss falls rapidly within the first 35 iterations, and begins to converge gradually in about 60 epochs. Figure 6 (b) shows that the model achieves earlier convergence within the first 30 iterations.
Figure 6: Convergence of FANS on datasets CIFAR10 and MNIST with sample sizes $|\mathcal{E}|=20,|\mathcal{E}|=200$ and $|\mathcal{E}|=2,000$.
6 Related Work
Feature attribution. Explainable artificial intelligence (XAI) methods aim to provide human-readable explanations to help users comprehend and trust the outputs created by ML models (Zhang & Zhu, 2018). At present, a series of XAI methods have been developed, mainly including building interpretable models (e.g., linear regression model) (Chen et al., 2023) and generating post hoc explanations. Feature attribution methods (FAMs) (Nielsen et al., 2022; Yang et al., 2023; Deng et al., 2024), which provide an estimate of the contribution of input features to a model’s prediction, is one of the most popular post hoc techniques. In this paper, we restrict our attention to feature attribution. Existing FAMs mainly measure the contribution of features to prediction through a perturbation test, i.e., by perturbing the features and comparing the resulting differences in predictions. Typically, Shapley value-based methods (Shapley et al., 1953; Lundberg & Lee, 2017; Zhang et al., 2021; Ren et al., 2021) measure the difference in a model’s predictions given data points with and without the feature included. Gradient-based methods (Sundararajan et al., 2017; Yang et al., 2023) measure the rate of change in prediction corresponding to a tiny change in an input feature. LIME-based methods (Ribeiro et al., 2016; Dhurandhar et al., 2022) learn simplified, interpretable models on perturbed feature subsets. In another line, subset-wise FAMs (e.g., GNNExplainer (Ying et al., 2019)) attempt to find a feature subset with the highest contribution by optimizing a loss function consisting of the $L_{p}$ norm of a feature mask and the prediction given the masked inputs (Fong & Vedaldi, 2017; Wang et al., 2021; Tan et al., 2022; Lucic et al., 2022; Zhang et al., 2023). However, current FAMs may not accurately distinguish the contributions between different features when their change in prediction is the same after perturbation. To address this issue, we extend feature attribution to compare the differences of each feature in PNS.
Probability of Necessity and Sufficiency (PNS). Necessity and sufficiency are two different aspects of causation. Sufficiency assesses the probability of a cause actively generating an effect, whereas necessity evaluates whether the cause, when absent, is capable of altering the effect, thereby measuring the effect’s dependence on the cause. At present, the main methods for feature attribution that consider sufficiency or necessity are LEWIS (Galhotra et al., 2021), LENS (Watson et al., 2021), and NSEG (Cai et al., 2022). However, since LEWIS and LENS only use PS or PN, the contribution score of an important feature subset may remain unchanged after including redundant features or removing some features inside. For example, we know that ‘fuel’, ‘oxygen’ and ‘heat’ cause ‘burning’. Since these three key factors and the ‘ice cream sold out’ (redundant feature) can also cause burning, we have PS(fuel, oxygen, heat) = PS(fuel, oxygen, heat, ice cream sold out). Another example is that without ‘oxygen’, ‘burning’ doesn’t happen. Thus, PN(fuel, oxygen, heat) = PN(oxygen). NSEG extracts the feature subset with the largest PNS by introducing an identifiability assumption and then optimizing the lower bound of PNS, but our FANS does not require this assumption since we directly compute PNS according to its definition, hinging on the outcome of $\mathbf{Y}$ after the intervention can be observed in our causal model.
@@ 7 Conclusion
This paper presents a novel attribution method called Feature Attribution with Necessity and Sufficiency (FANS) that can better distinguish the contribution of different feature subsets to predictions. We first formally define the Structural Causality Model (SCM) for the perturbation test, which is widely used by most existing feature attribution methods. Building upon this SCM, FANS introduces Necessary and Sufficient Attribution, defined as the highest probability in the Probabilities of being a Necessity and Sufficiency (PNS) cause of the prediction change for perturbing samples in different neighborhoods, through our proposed heuristic strategy for estimating the neighborhood and dual-stage (factual and interventional) perturbation test. To generate counterfactual samples, we propose to adopt a sampling-importance-resampling approach. Moreover, we show the superior performance of our FANS compared to state-of-the-art methods. The main limitation of our method is that the estimation methods of the boundary $b$ of the target input’s neighborhood and threshold $c$ for prediction change are both heuristic. In future work, it would be meaningful to provide a theoretical framework on the identifiability of $b$ and $c$.
Impact Statement
@@
This paper proposes a feature attribution method called FANS, which can distinguish the contributions of different features even if the results change similarly after perturbation. FANS can be used in medical diagnosis, autonomous driving, and other fields to assist people in understanding the predictions of the model, which have the potential to facilitate the application of existing artificial intelligence technologies in real-world scenarios.
Acknowledgements
This research was supported in part by National Key R&D Program of China (2021ZD0111501), National Science Fund for Excellent Young Scholars (62122022), and Natural Science Foundation of China (62206064, 62206061), and was supported by Turing AI Fellowship under grant EP/V023756/1. We thank Wenlin Chen for valuable suggestions, and the anonymous reviewers who gave useful comments.
Appendix A Evaluation of PN and PS
As illustrated in Figure 1, the causal diagram demonstrates that variable $\mathbf{Y}$ is influenced by variables $\tilde{\mathbf{X}}$ and $\mathbf{S}$, with a function of $\mathbf{Y}=f(g(\tilde{\mathbf{X}},\mathbf{S},\mathbf{x}^{\prime}))$ (Eq. 2). When variables $\tilde{\mathbf{X}}$ and $\mathbf{S}$ are intervened as $\mathbf{x}$ and $\bar{\mathbf{s}}$ respectively, according to the counterfactual reasoning paradigm (Pearl, 2022), we have:
$P(\bar{B}_{\mathbf{s},c_{\bar{A}_{\mathbf{s},b}}}\mid A_{\mathbf{s},b},B_{\mathbf{s},c})=\mathbb{E}_{\mathbf{x}\sim P(\tilde{\mathbf{X}}\mid A_{\mathbf{s},b},B_{\mathbf{s},c})}\left[P\left(\left|\mathbf{Y}-f(\mathbf{x})\right|\leq c\right)_{\operatorname{do}(\tilde{\mathbf{X}}=\mathbf{x},\mathbf{S}=\bar{\mathbf{s}})}\right]$
For simplicity, we replace $\mathbf{Y}$ with its value $f(g(\tilde{\mathbf{X}}=\mathbf{x},\mathbf{S}=\bar{\mathbf{s}},\mathbf{x}^{\prime}))$, and since $\mathbf{S}$ is an exogenous variable, the above equation can be written as
$=\mathbb{E}_{\mathbf{x}\sim P(\tilde{\mathbf{X}}\mid A_{\mathbf{s},b},B_{\mathbf{s},c})}\left[P\left(\left|f\left(g\left(\mathbf{x},\bar{\mathbf{s}},\mathbf{x}^{\prime}\right)\right)-f(\mathbf{x})\right|\leq c\right)_{\operatorname{do}(\tilde{\mathbf{X}}=\mathbf{x})}\right]$
Similarly, the derivation of Eq. 9 is given by
$P\left(B_{\mathbf{s},c_{A_{\mathbf{s},b}}}\mid\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c}\right)$ $=\mathbb{E}_{\mathbf{x}\sim P\left(\tilde{\mathbf{X}}\mid\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c}\right)}[P\left(\left|\mathbf{Y}-f(\mathbf{x})\right|>c\right)_{\mathrm{do}(\tilde{\mathbf{X}}=\mathbf{x},\mathbf{S}=\mathbf{s})}]$
$=\mathbb{E}_{\mathbf{x}\sim P(\tilde{\mathbf{X}}\mid\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})}[P\left(\left|f\left(g\left(\mathbf{x},\mathbf{s},\mathbf{x}^{\prime}\right)\right)-f(\mathbf{x})\right|>c\right)_{\mathrm{do}(\tilde{\mathbf{X}}=\mathbf{x})}]$
Appendix B Resampling Weight for Approximate Sampling
@@
SIR consists of three steps. 1) Draw samples $\mathbf{x}_{1}$, …, $\mathbf{x}_{k}$ from the proposal distribution $Q(\mathbf{X})$. 2) For each $\mathbf{x}_{i}$, calculate the weight $w(\mathbf{x}_{i})=\frac{P(\mathbf{x}_{i}|\bar{A}_{\mathbf{s},b},\bar{B}_{\mathbf{s},c})}{Q(\mathbf{x}_{i})}$ if $\|\mathbf{x}_{\bar{\mathbf{s}}}-\mathbf{x}^{t}_{\bar{\mathbf{s}}}\|_{p}\leq b$ (definition of variable $\tilde{\mathbf{X}}$ in Eq. 1) and 00 otherwise. 3) Draw samples from $\mathbf{x}_{1}$, …, $\mathbf{x}_{k}$ based on their weights.
[…]
Table 3: Notation and Descriptions.
Notations Descriptions
$f$ and $g$ The black box model and perturbation function.
$\mathbf{X}$ The variable of the input.
$\mathbf{x}$ A particular value of $\mathbf{X}$.
$\mathbf{x}^{\prime}$ A baseline for perturbation.
$\mathbf{x}^{t}$, Target input to be explained
$\tilde{\mathbf{X}}$ Input with fixed features on $\mathbf{S}$ that are similar to $\mathbf{x}^{t}$.
$\mathbf{Y}$ The variable of the model’s prediction after perturbation.
$\mathbf{y}$ A particular value of $\mathbf{Y}$.
$\mathbf{m}$ A random mask used to perturb the input.
$\mathbf{S}$ A set containing the dimensions of $\mathbf{X}$ specified to be perturbed. $\mathbf{S}$ can be any set and is not fixed.
$\mathbf{s}$ A particular example of $\mathbf{S}$, indicating the dimensions of the actual perturbation.
$\mathbf{x}_{\mathbf{s}}$, $\mathbf{m}_{\mathbf{s}}$ Subvectors of $\mathbf{x}$ and $\mathbf{m}$ that are indexed by $\mathbf{s}$.
$b$ Boundary of the neighborhood of the target input (Eq. 7).
$c$ Threshold used to determine if there is a significant difference between the perturbed and
original predictions.
$A_{\mathbf{s},b}$, $\bar{A}_{\mathbf{s},b}$ Perturbation event and its complement.
$B_{\mathbf{s},c}$, $\bar{B}_{\mathbf{s},c}$ Prediction event and its complement.
$r$, $d$ Normalization constant and the dimensionality of the input.
$t$, $u$ Number of perturbations in the sufficiency and necessity modules, respectively.
$w_{\text{SF}}$ and $w_{\text{NC}}$ Sample weighting functions for PS and PN, respectively.
$\mathcal{E}_{\text{SF}}$ and $\mathcal{E}_{\text{NC}}$ Sample sets drawn from conditional distributions of PS and PN through SIR, respectively.
$\mathbb{I}(\cdot)$ and $\circ$ Indicator function and element-wise multiplication.
Appendix C Experimental Setup
C.1 Datasets and Target Models